COMPLETE STRUCTURE OF THE ANTI ROTAMER OF 1,1,2,2-TETRAFLUOROETHANE FROM IN-FRARED SPECTROSCOPY

NORMAN C. CRAIG, CATHERINE M. OERTEL, and \underline{DAVID C. OERTEL}, Department of Chemistry, Oberlin College, Oberlin, OH 44074; MICHAEL LOCK, Physikalisch-Chemisches Institut, Justus-Liebig-Universität, D-35392 Giessen, Germany.

High-resolution (0.0018 cm^{−1}) infrared spectra have been recorded in the CF stretching region for 1,1,2,2-tetrafluoroethane and its d_{2} and ^{13}C_{2} isotopomers. These bands are sufficiently intense to give useful gas-phase spectra at −100^{∘}C in a 3-m cell. For each spectrum the rotational structure of the A/C-type and B-type bands due to the anti rotamer, which is a highly asymmetric rotor with κ = -0.295, have been analyzed. Rotational constants for a Watson-type Hamiltonian have been fitted for the common ground state of each isotopomer and the separate upper states. A complete structure for the nonpolar anti rotamer has been derived and compared with the structure of the polar gauche rotamer obtained from microwave spectroscopy. The geometric parameters for the two rotamers are compared with recent calculations, and trends in these parameters are discussed.